# RobMOT: Robust 3D Multi-Object Tracking by Observational Noise and State Estimation Drift Mitigation on LiDAR PointCloud

### v2

@@ Abstract
    We propose an online 3D MOT framework with exceptional tracking performance, surpassing recent benchmarks to show its robustness across different datasets and detectors. The tracking performance margin of our framework with the state-of-the-art CasTrack [1] reaches $4\%$ in HOTA and $6\%$ in MOTA, and the enhancement in computational latency is up to $53\%$, as demonstrated in Figure 6.
@@ 3.1 Evaluation with Benchmarks
3.1.1 Quantitative Evaluation
Table 1: Tracking performance on KITTI dataset testing split with the latest MOT benchmarks
Method HOTA MOTA AssA AssRe IDSW CasTrack* [1] 81% 91.9% 84.2% 87.6% 24 PC-TCNN [26] 80.9% 91.7% 84.1% 87.5% 37 Rethink MOT [5] 80.1% 91.5% 83.6% 87.6% 46 MSA-MOT [27] 78.5% 88.0% 82.6% 85.2% 91 DeepFusion-MOT [2] 75.5% 84.6% 80.1% 82.6% 84 PC3T [28] 77.8% 88.8% 81.6% 84.8% 225 Mono-3D-KF [29] 75.5% 88.5% 77.6% 80.2% 162 PolarMOT [30] 75.2% 85.1% 76.95% 80.0% 462 TripletTrack [31] 73.58% 84.32% 74.66% 77.3% 322 RobMOT* (Ours) 81.22% 90.5% 85.8% 89.7% 7
    Methods marked with * use the same detector
Initially, we evaluate the tracking performance on the KITTI testing dataset with the latest state-of-the-art tracking methods listed in the KITTI leaderboard. As illustrated in Table 1, our framework shows a dramatic drop in the ID switch metric (IDSW) compared to the benchmarks, with a difference of 17 IDs, which shows the tracking stability of the proposed approach. Additionally, the framework outperforms in the Higher Order Tracking Accuracy (HOTA) and association accuracy (AssA) metrics with margins of $0.22\%$ and $1.6\%$, respectively. However, CasTrack is superior in Multi-object Tracking Accuracy (MOTA) metrics with a margin of $1.4\%$. Although HOTA measures the consistency of the tracking performance by combining association, localization, and detection compared to MOTA [32], Table 1 shows comparable performance between the proposed solution and CasTrack [1] in HOTA and MOTA metrics despite the frequent change in ID during tracking for CasTrack that reaches $24$ times.
Another evaluation is conducted on WOD testing dataset. The dataset consists of simple to moderate tracking conditions labeled “Level 1" and challenging conditions labeled “Level 2" in Table 4. Our proposed solution shows superior tracking performance on the recent benchmark, including CasTrack [1], with a reduction in the false positives ratio (FP).
Table 2: Tracking performance on Waymo Open dataset testing split with the latest MOT benchmarks
@@ 3.1.1 Quantitative Evaluation
    Methods marked with * use the same detector
Despite the superior performance that our approach achieves in KITTI (Table 1) and WOD (Table 2) datasets, CasTrack [1] shows close tracking performance. Thus, we perform a sensitive and qualitative analysis with CasTrack [1] to show the capability of the two methods under different detectors and challenging tracking conditions.
The performance of tracking-by-detection approaches is highly dependent on the utilized detector. High detection performance enhances tracking performance, making evaluating a tracking approach a tricky task. To ensure the superiority of a tracking method, it should maintain its outperformance across various detectors that confirm the contribution of the tracking method. To this end, we conduct a sensitivity analysis across five detectors with different detection performances on KITTI evaluation dataset and compare the proposed approach with CasTrack [1]. Figure 6 shows the tracking performance across the detectors using three different metrics: MOTA, HOTA, and IDSW. Although both methods show close performance in HOTA and MOTA with VirConv [19], CasA [20], and PointRCNN [22], CasTrack [1] performance drops dramatically in PV-RCNN [23] with a margin of more than $6\%$ in MOTA and $4\%$ in HOTA. Meanwhile, our proposed method shows superiority across almost all detectors while maintaining the lowest IDSW compared to CasTrack [1]. These results validate the contribution of the proposed approach while maintaining superior performance across different detectors. In addition, the proposed approach achieved higher frame per second (fps) than CasTrack [1] across the detectors, with enhancement in fps reaching $+59\%$ with the Second [21] detector.

### v3

@@ Abstract
    We propose an online 3D MOT framework with exceptional tracking performance, surpassing recent benchmarks to show its robustness across different datasets and detectors. The tracking performance margin of our framework with the state-of-the-art CasTrack [1] reaches $4\%$ in HOTA and $6\%$ in MOTA, and the enhancement in computational latency is up to $53\%$, as demonstrated in Figure LABEL:fig:hota_mota_comp.
@@ 3.1 Evaluation with Benchmarks
3.1.1 Quantitative Evaluation
Table 1: The table compares the latest benchmarks listed in the KITTI leaderboard on the testing split. The highest performance at each metric is highlighted in red, the second highest in blue, and the third highest in green. Methods that use the CasA [20] detector are marked by (*), while ($\dot{+}$) is used for methods that use the VirConv [19] detector.
Method HOTA MOTA AssA AssRe IDSW TripletTrack [26] 73.58% 84.32% 74.66% 77.3% 322 PolarMOT [27] 75.2% 85.1% 76.95% 80.0% 462 DeepFusion-MOT [2] 75.5% 84.6% 80.1% 82.6% 84 Mono-3D-KF [28] 75.5% 88.5% 77.6% 80.2% 162 PC3T [29] 77.8% 88.8% 81.6% 84.8% 225 MSA-MOT [30] 78.5% 88.0% 82.6% 85.2% 91 UG3DMOT* [31] 78.6% 87.98% 82.28% 85.36% 30 CasTrack* [1] 80% 90.5% 82.7% 86.4% 150 Rethink MOT [5] 80.1% 91.5% 83.6% 87.6% 46 MCTrack [32] 80.78% 89.82% 84.29% - 65 PC-TCNN [33] 80.9% 91.7% 84.1% 87.5% 37 $\dot{+}$ [19, 1] 81.56% 90.54% 85.2% 88.7% 113 RobMOT* (Ours) 81.22% 90.5% 85.8% 89.7% 7 $\dot{+}$ (Ours) 81.76% 91.02% 85.6% 89.25% 7
We evaluate the performance of the proposed framework, RobMOT [nagy2024robmotrobust3dmultiobject], against several state-of-the-art multi-object tracking benchmarks on the KITTI dataset testing split. The results are summarized in Table 1, which presents metrics: HOTA (Higher Order Tracking Accuracy), MOTA (Multi-Object Tracking Accuracy), AssA (Association Accuracy), AssRe (Association Recall), and IDSW (Identity Switches).
In this comparison, we use two different detectors, VirConv [19] and CasA [20], to evaluate the tracking performance rather than the detection fairly. We are re-running the official implementation of the CasTrack [1] work as it is the latest top-performing classical approach using their proposed detectors CasA [20] and VirConv [19]. Our proposed framework, RobMOT, outperforms CasTrack with the CasA detector, marked by (*) in Table 1, with about 1.22% in HOTA. By switching the detector to VirConv, our method still outperforms CasTrack [1], named "VirConvTrack" [19] in Table 1, with about 0.2% in HOTA, marked by ($\dot{+}$) in Table 1. These results show the contribution of the proposed tracking approach regardless of the performance of the utilized detector. Notably, RobMOT is ranked the first or second highest in most metrics, with CasA or VirConv detector among the other benchmarks.In addition, the proposed work excels in Association Accuracy (AssA) with 85.8%, marking it the highest among the evaluated methods. Similarly, it achieves a leading Association Recall (AssRe) of 89.7%, indicating its robust capability to maintain correct identity associations throughout the tracking process.
One of the most significant contribution is the low Identity Switches (IDSW) score of 7 for RobMOT with CasA and VirConv detectors. This demonstrates the model’s ability to consistently track object identities over time, which is critical for reliable multi-object tracking. In contrast, other benchmark methods, such as UG3DMOT [31] and CasTrack [1] with the same detector, exhibit substantially higher IDSW scores of 30 and 150, respectively, indicating more frequent identity switches.
@@ 3.1.1 Quantitative Evaluation
Our further experiments on the KITTI evaluation dataset across five detectors show CasTrack [1] and VirConvTrack [19] dependency on the utilized detector, as illustrated in Figure 6. It shows the tracking performance across the detectors using three different metrics: MOTA, HOTA, and IDSW. Although both methods show close performance in HOTA and MOTA with VirConv [19], CasA [20], and PointRCNN [22], CasTrack [1] performance drops dramatically in PV-RCNN [23] with a margin of more than $6\%$ in MOTA and $4\%$ in HOTA.
Our proposed method, while maintaining the lowest IDSW compared to CasTrack [1], consistently outperforms across all detectors. These results validate the contribution of the proposed tracking approach, demonstrating its exceptional performance across various detectors. Additionally, the proposed approach achieves higher frame per second (fps) than CasTrack [1] across the detectors, with an fps enhancement reaching $+59\%$ with the Second [21] detector.
Another evaluation is conducted on the WOD testing dataset to measure the generalization of the approach. The dataset consists of simple to moderate tracking conditions labeled “Level 1" and challenging conditions labeled “Level 2" in Table 2. Our proposed solution shows superior tracking performance on the recent benchmark, including CasTrack [1], with a reduction in the false positives ratio (FP).
Table 2: Tracking performance on Waymo Open dataset testing split with the latest MOT benchmarks. Methods that use the CasA [20] detector are marked by (*)
@@
In conclusion, the proposed approach excels in tracking stability and accuracy, significantly reducing ID switches and enhancing fps performance across various detectors. It outperforms recent state-of-the-art methods like CasTrack [1] on both the KITTI and WOD datasets, demonstrating robust generalization and efficiency. The quantitative results validate the contribution and exceptional performance of the proposed framework.